OpenStudy (anonymous):

Find the 6th term of the sequence with t1 = -4 and tn = 5tn-1.

OpenStudy (campbell_st):

is the equation \[t_{n} = 5t_{n -1}\]

OpenStudy (anonymous):

yeah

OpenStudy (campbell_st):

ok... so its basically 5 x the previous term so \[t_{2} = 5\times t_{2 -1}...... t_{2} = 5 \times t_{1} \] you can repeat the process... until you get to the 6th term or you can use the formula \[T_{n} = ar^{n -1} \] to find a term in a geometric series.. a = 1st term r = common ratio... n = number of terms

OpenStudy (anonymous):

ok thanks